OpenStudy (anonymous):

8d - 1/3(6 - 9d) = 42 Solve for d

OpenStudy (noorik.):

8d-1 = 42*(6-9d) 8d-1 = 252-378d 386d=253 d= 386/253

jhonyy9 (jhonyy9):

so i think that this wan being 8d - (6-9d)/3 =42 8d - 3(2-3d)/3 =42 8d -2 +3d = 42 11d = 44 d = 44/4 d = 11